b) A first order reactionA→B, with the rate constant k is taking place in CSTR fed with A at concentration Car which remains unchanged. There are likely to be some deviations in the feed rate (F) of A. Derive linearized transfer function between. concentration of A in the outlet and feed rate of A assuming that volume V of reacting mixture remains unchanged.


1
Expert's answer
2021-09-21T02:06:58-0400


 Is the model linear? If we decide to solve the model numerically, we do not have to linearize; in fact, the non-linear model would be more accurate. However, in this problem we seek the insight obtained from the approximate, linear model.

All terms involve a constant times a variable (linear) except for the following term, which is linearized using the Taylor series..

"(1) C0.5 \u2248 (C0.5 ) + 0.5(C\u22120.5 ) (C \u2212 C )"

higher order terms AAs AsAAs

This approximation can be substituted into equation 2, and the initial steady-state model subtracted to obtain the following, with C’A = CA - CAS.

"(2) V dC'A = F(C' \u2212FC' ) \u2212 Vk(0.5C\u22120.5 )C' dt" A0

This linear, first order ordinary differential equation model can be arranged into the standard form, given in the following.


"\u03c4(sC'A (s) \u2212 C'A (t) |t=0 )+ C'A (s) = KC'A0 (s)"